OpenStudy (anonymous):

x+1/x+2=6 is given. Find (x+2)^2+1/(x+2)^2.

OpenStudy (danjs):

\[\frac{ x+1 }{ x+2 } = 6\]

OpenStudy (danjs):

that right?

OpenStudy (danjs):

(x+1)/(x+2) = 6

OpenStudy (anonymous):

\[x+\frac{ 1 }{x+2 }=6, (x+2)^2+\frac{ 1 }{ (x+2)^{2}}=?\]

OpenStudy (danjs):

solve the equation for x, then use that x in the expression to find a value

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & x+2+\frac{ 1 }{x+2 }=8, \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & put\ \ x+2=a \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

ok thank you i could not think i should add 2 to both sides so the answer is 62

OpenStudy (mathmath333):

ans square both sides

ganeshie8 (ganeshie8):

that is clever! \(a+\frac{1}{a} = k \implies a^2+\frac{1}{a^2} = k^2-2\)

OpenStudy (danjs):

yeah, looks like that was the point of the problem.. nice